# Twist-free axisymmetric critical collapse of a complex scalar field

### v1

@@ I Introduction
After thirty years of progress in the field, there is very strong evidence that the Choptuik critical solution for the scalar field model is universal in spherical symmetry. It has been systematically validated in the ‘numerical laboratory’ of various independent gravitational collapse codes, with various formulations of GR and discretizations. Nevertheless, the extent to which this universality carries over with initial data containing arbitrary non-spherical perturbations is not absolutely clear. In fact, numerical work shows that the Choptuik solution is not universal under a more relaxed symmetry hypothesis. More generally, evidence suggests that the threshold of collapse becomes more subtle as symmetry is discarded.
@@
In fact, the vacuum scenario differs qualitatively from Choptuik’s classical setup, because it does not admit a non-trivial dynamical spherical limit in 3+1 dimensions. This makes the understanding of how it relates to the standard picture of critical collapse tricky. The urge to interpret subtleties appearing in the axisymmetric category leads naturally to the thought of studying models that admit dynamical black hole formation in spherical symmetry and with arbitrarily small axisymmetric deformations. For the real scalar field such configurations have been studied in a number of works. Nonlinear numerical evolutions were presented by Choptuik et al. [15] and Baumgarte [16] in axisymmetry and by Healy and Laguna in [17], and Deppe et al. [18], in full 3d. In [19], Clough and Lim study the critical collapse of a real scalar field, but with a nonlinear potential. Unsurprisingly all of the studies in full 3d without symmetry achieve less tuning to the threshold of collapse. Again in full 3d, but working perturbatively to linear order, Martín-García & Gundlach [20] found that all nonspherical mode solutions on the Choptuik spacetime decay. In [16] an interpretation was offered that with large enough departures from sphericity the slowest of these decay rates actually flips sign such that aspherical modes grow. Thus, consistent with the results of [15], Baumgarte finds that this growth leads to a bifurcation of the data in which two centers of collapse appear.
@@ II Massless complex scalar field
Writing $\psi=\psi^{\textrm{Re}}+i\psi^{\textrm{Im}}$, the stress energy tensor can be decomposed as $T_{\mu\nu}=T_{\mu\nu}^{\textrm{Re}}+T_{\mu\nu}^{\textrm{Im}}$, with each of these corresponding to the standard expression for the stress-energy tensor of a real massless scalar field, evaluated on the real and imaginary part of the field, respectively. Such a split will not, in general, be possible in the presence of a potential, but does make the interpretation of the dynamics straightforward in the massless case. Due to the fact that the real and imaginary parts of the complex scalar field follow exactly the same equation as the real scalar field itself, interacting only through their minimal coupling to gravity, we expect a qualitative behavior similar to the real scalar field case. In this way, one could, in principle, consider the massless complex scalar field case as a supplementary study of the real scalar field case, albeit with a bigger solution space and correspondingly richer dynamics due to the global internal $U(1)$ symmetry of the complex field.
@@ III.1.1 Geometry
where $\alpha\gamma_{0}=\alpha\gamma_{2}=2$, $\gamma_{1}=-1$, and $\gamma_{4}=\gamma_{5}=\frac{1}{2}$. We define the lapse $\alpha$, shift $\beta^{i}$ and spatial metric $\gamma_{ij}$ according to the standard notation in NR. Following the suggestion presented in [26], we have adapted our damping parameters $\gamma_{0}$ and $\gamma_{2}$ to collapse spacetimes by adding the lapse in their definition. At the outer boundary we employ constraint-preserving, radiation-controlling boundary conditions, see [30, 31, 24] for details.
@@ IV Phase space search
Hence, the task of locating an apparent horizon simply transforms to a root finding process and the threshold search is performed via a binary search algorithm. In the case of axisymmetry, however, one has to solve a differential equation instead. The numerical tool we are using is AHloc3d [39], which uses a flow method starting from a large arbitrary 2d surface, in combination with a Newton-Raphson method in the end. A shortcoming of this method is that for a horizon that is not star-shaped our tool is unable to locate it, since this assumption is used explicitly in the search algorithm. Moreover, it is a post-processing tool that requires expensive output, both in terms of time and storage, from our evolution code. Our search strategy at each bisection level is thus to perform a series of runs, with some of them obviously dispersing to infinity (subcritical) and some others blowing up (potentially supercritical). Subsequently, we reproduce the ones that blow up, starting from check-pointed data late in the evolution, but now additionally writing output that is destined as an input for AHloc3d which then provides a proper classification. For a detailed review of horizon searches in NR see [40]. For more recent reviews in the form of PhD theses, see [41, 42].
@@ V.1 Spherical Symmetry
The initial conditions (amplitude $a^{\textrm{Re/Im}}$, location of center $r_{0}^{\textrm{Re/Im}}$ and width $s_{z}^{\textrm{Re/Im}}=s_{\rho}^{\textrm{Re/Im}}\equiv s^{\textrm{Re/Im}}$) for the Gaussian pulses of our spherical families are shown in Table 2. Tuning either the real or the imaginary part is equivalent, since they take part in the equations of motion identically. To demonstrate this, we are alternating between tuning the real or imaginary part in different families. We have evolved all families using two different gauges, first that appearing in [18] and also that of [26], which we refer to as HDWG-$\ln(\alpha)$ and DSS-compatible respectively. We only present results from the gauge choice that gives the cleanest results for each family. One of the families is ingoing, denoted as “ing”, whilst the rest start at a moment of time symmetry, denoted as “ts”.
@@
fitting the values by regression. We find values around $\gamma\simeq 0.37$ for all families. This coincides with the familiar value from simulations of the real scalar field in spherical symmetry.
@@ V.2 Departures from Spherical Symmetry
Numerical data produced with bamps has been thoroughly convergence tested in numerous challenging validation runs, see for instance [24, 25, 32]. Doing so carefully is a technically tricky task with any pseudospectral method, because one cannot expect a simple power-law improvement in errors with resolution as in a finite-differencing method. Instead, in the best case scenario one should obtain exponential convergence with resolution, but typically with an unknown exponent. (To avoid misinterpretation: exponential convergence is highly desirable, but complicates quality evaluation). These difficulties are amplified yet further when using mesh-refinement, because as resolution is increased, the grid hierarchies will only match if carefully specified at the outset, which requires more book-keeping to set up. For applications in critical collapse, the situation is yet worse because, by design, we consider spacetimes which depend sensitively on the chosen initial data. Consequently numerical error can have an outsized effect on the outcome of the evolution and can, for instance, even kick the end-state of the evolution from dispersion to black hole formation, and thus end up with quite different levels of constraint violation. For these reasons, and the fact that to do so would be computationally highly expensive, we have not performed systematic convergence tests on all of the data presented here. In particular, our mesh-refinement infrastructure was tested in detail in [25], including with pre-prescribed grid hierarchy. In [32] detailed validation runs were presented with exactly the complex scalar field code we employ presently. For this project specifically we performed such tests for strong data far from the threshold of collapse and saw convincing convergence with resolution. Confidence in the quality of our aspherical near-threshold evolutions is bolstered by a number of facts. First, we monitor constraint violation throughout. The constraints do grow as the dynamics become increasingly violent at the threshold of collapse, but they remain at least five orders of magnitude, and often more, smaller than any curvature scalars with equivalent units that we consider. Second, we have already compared above our spherical runs with the (admittedly limited) available literature for the complex scalar field, and find excellent agreement. Below we will reveal that our results agree qualitatively with published work for the analogous setup for the real scalar field. All these points considered, we are therefore confident in the qualitative behavior we find, especially when considering small asphericities, but do not feel justified in stating strict error bounds, which are left for future work.
@@
For instance, it is possible for the bifurcation to occur such that subsequent curvature peaks occur ever closer to the origin so that in the limit of infinite tuning, the curvature singularity forms at the origin. On the other hand, curvature singularities could form away from the origin. In either case, the blow-up could exhibit DSS or not. At the level of tuning currently possible in vacuum collapse, evidence suggests that there are families of initial data in which the singularity forms at the origin and others in which it does not. In either case, there is no evidence for strict DSS as in the spherical scalar field setting. But there are families that exhibit approximate DSS. Since the dynamics of the complex scalar field necessarily include those of the vacuum and real scalar field cases, all of this is highly relevant. Yet another possibility, which we do not presently have the tools to diagnose, is the formation of a naked singularity on an open set of parameters within a family.
@@
Taken together, Figs. 7 and 9 indicate that near the threshold, our most aspherical family does not exhibit discrete self-similarity centered at the origin and is thus not a single copy of the Choptuik solution. That said, we could still reasonably describe a portion of the spacetime as approximately DSS about the center. They also show that the gravitational wave component of the spacetime becomes much more significant in the highly aspherical regime. Figure 8 corroborates this, and furthermore suggests that a lack of DSS may be more clearly diagnosed by looking at a more complete set of scalar fields.
@@ VI Conclusions
Our first set of findings concerns spherical symmetry configurations, for which we tuned four distinct families of initial data to the threshold of collapse. Although closely related models have been studied before, as far as we are aware, this is the first time that the spherical configuration for the massless complex field has been carefully treated. We found strong evidence that there is a universal critical spacetime metric which agrees with that of the real scalar field. This metric is characterized partially by the echoing period $\Delta\simeq 3.4$. Perturbations from the critical solution lead to power-law growth of scalar quantities with exponent $\gamma\simeq 0.37$ also familiar from the spherical real scalar field. Since the complex scalar solution space includes that of the real field, this was certainly a possibility a priori, but requires the real and imaginary parts of the field to concoct precisely the stress-energy tensor of the Choptuik solution, and so was a surprise to us. Interestingly similar behavior in which, in a certain limit, the dynamics of a more complicated system agree with those of a simpler model, occurs at the threshold of blow-up for several of the models examined in [50]. This surprise is heightened by the fact that, in the general axisymmetric setting with twist (angular momentum) [21] the threshold of collapse for real and complex fields certainly do not coincide; in that context the real field cannot support angular momentum whereas the complex field can.
@@ Acknowledgments
We acknowledge financial support provided by FCT/Portugal through grants 2022.01324.PTDC, PTDC/FIS-AST/7002/2020, UIDB/00099/2020 and UIDB/04459/2020. D.C. acknowledges support from the by the STFC Research Grant No. ST/V005669/1 "Probing Fundamental Physics with Gravitational-Wave Observations". D.C. and F.A. were supported in part by the Deutsche Forschungsgemeinschaft (DFG) under Grant No. 406116891 within the Research Training Group RTG 2522/1. H.R.R acknowledges support from the Fundação para a Ciência e Tecnologia (FCT) within the projects UID/04564/2021, UIDB/04564/2020, UIDP/04564/2020 and EXPL/FIS-AST/0735/2021 and the H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. The figures in this article were produced with Plots.jl [67], Matplotlib [68], A. Khirnov’s nr_analysis_axi package [69] and ParaView [70, 71]. Several calculations were performed using the xAct package [72] for Mathematica. Numerical simulations were performed at the Leibniz Supercomputing Centre (LRZ), supported by the project pn36je. The authors thankfully acknowledge the computer resources, technical expertise and assistance provided by CENTRA/IST. Computations were performed at the cluster “Baltasar-Sete-Sóis” and supported by the H2020 ERC Consolidator Grant "Matter and strong field gravity: New frontiers in Einstein’s theory" grant agreement no. MaGRaTh-646597.

### v2

@@ I Introduction
After thirty years of progress in the field, there is very strong evidence that the Choptuik critical solution for the scalar field model is universal in spherical symmetry. It has been systematically validated in the ‘numerical laboratory’ of various independent gravitational collapse codes, with various formulations of GR and discretizations [2]. Nevertheless, the extent to which this universality carries over with initial data containing arbitrary non-spherical perturbations is not absolutely clear. In fact, numerical work shows that the Choptuik solution is not universal under a more relaxed symmetry hypothesis. As we discuss in detail momentarily, evidence suggests that the threshold of collapse becomes more subtle as symmetry is discarded.
@@
In fact, the vacuum scenario differs qualitatively from Choptuik’s classical setup, because it does not admit a non-trivial dynamical spherical limit in 3+1 dimensions. This makes the understanding of how it relates to the standard picture of critical collapse tricky. The urge to interpret subtleties appearing in the axisymmetric category leads naturally to the thought of studying models that admit dynamical black hole formation in spherical symmetry and with arbitrarily small axisymmetric deformations. For the real scalar field such configurations have been studied in a number of works. Nonlinear numerical evolutions were presented by Choptuik et al. [15], Baumgarte [16] and Gundlach et al. [17] in axisymmetry and by Healy and Laguna in [18], and Deppe et al. [19], in full 3d. In [20], Clough and Lim study the critical collapse of a real scalar field, but with a nonlinear potential. Unsurprisingly all of the studies in full 3d without symmetry achieve less tuning to the threshold of collapse. Again in full 3d, but working perturbatively to linear order, Martín-García & Gundlach [21] found that all nonspherical mode solutions on the Choptuik spacetime decay. In [16] an interpretation was offered that with large enough departures from sphericity the slowest of these decay rates actually flips sign such that aspherical modes grow. Thus, consistent with the results of [15], Baumgarte finds that this growth leads to a bifurcation of the data in which two centers of collapse appear.
@@ II Massless complex scalar field
Writing $\psi=\psi^{\textrm{Re}}+i\psi^{\textrm{Im}}$, the stress energy tensor can be decomposed as $T_{\mu\nu}=T_{\mu\nu}^{\textrm{Re}}+T_{\mu\nu}^{\textrm{Im}}$, with each of these corresponding to the standard expression for the stress-energy tensor of a real massless scalar field, evaluated on the real and imaginary part of the field, respectively. Note that since there is no direct interaction between the real and imaginary parts of the complex field, so each individually satisfies the massless Klein-Gordon equation. Such a split will not, in general, be possible in the presence of a potential, but does make the interpretation of the dynamics straightforward in the massless case. Due to the fact that the real and imaginary parts of the complex scalar field follow exactly the same equation as the real scalar field itself, interacting only through their minimal coupling to gravity, we expect a qualitative behavior similar to the real scalar field case. In this way, one could, in principle, consider the massless complex scalar field case as a supplementary study of the real scalar field case, albeit with a bigger solution space and correspondingly richer dynamics due to the global internal $U(1)$ symmetry of the complex field.
@@ III.1.1 Geometry
where $\alpha\gamma_{0}=\alpha\gamma_{2}=2$, $\gamma_{1}=-1$, and $\gamma_{4}=\gamma_{5}=\frac{1}{2}$. We define the lapse $\alpha$, shift $\beta^{i}$ and spatial metric $\gamma_{ij}$ according to the standard notation in NR. Following the suggestion presented in [27], we have adapted our damping parameters $\gamma_{0}$ and $\gamma_{2}$ to collapse spacetimes by adding the lapse in their definition. At the outer boundary we employ constraint-preserving, radiation-controlling boundary conditions. These conditions work by setting the incoming gravitational wave degrees of freedom, captured by the Weyl scalar $\Psi_{0}$, to vanish, and likewise doing the same for all incoming characteristic variables of the constraint subsystem. See [31, 32, 25] for details.
@@ IV Phase space search
Hence, the task of locating an apparent horizon simply transforms to a root finding process and the threshold search is performed via a binary search algorithm. In the case of axisymmetry, however, one has to solve a differential equation instead. The numerical tool we are using is AHloc3d [40], which uses a flow method starting from a large arbitrary 2d surface, in combination with a Newton-Raphson method in the end. A shortcoming of this method is that for a horizon that is not star-shaped our tool is unable to locate it, since this assumption is used explicitly in the search algorithm. Moreover, it is a post-processing tool that requires expensive output, both in terms of time and storage, from our evolution code. Our search strategy at each bisection level is thus to perform a series of runs, with some of them obviously dispersing to infinity (subcritical) and some others blowing up (potentially supercritical). Subsequently, we reproduce the ones that blow up, starting from check-pointed data late in the evolution, but now additionally writing output that is destined as an input for AHloc3d which then provides a proper classification.
@@
We would prefer to have a more reliable apparent horizon finder, but would have to generalize our present algorithm to allow non star-shaped horizons to do so. For a detailed review of horizon searches in NR see [41]. For examples of non-star shaped horizon searches see [42, 43]. For more recent reviews in the form of PhD theses, see [44, 43]. Nevertheless, a better apparent horizon approach on the supercritical side would essentially result in a smaller value in our estimation of the threshold amplitude. Looking ahead, this would affect only figures across the phase space, as opposed to those from a fixed spacetime, exaggerating the results that we report below.
@@ V.1 Spherical Symmetry
The initial conditions (amplitude $a^{\textrm{Re/Im}}$, location of center $r_{0}^{\textrm{Re/Im}}$ and width $s_{z}^{\textrm{Re/Im}}=s_{\rho}^{\textrm{Re/Im}}\equiv s^{\textrm{Re/Im}}$) for the Gaussian pulses of our spherical families are shown in Table 2. Tuning either the real or the imaginary part is equivalent, since they take part in the equations of motion identically. To demonstrate this, we are alternating between tuning the real or imaginary part in different families. We have evolved all families using two different gauges, first that appearing in [19] and also that of [27], which we refer to as HDWG-$\ln(\alpha)$ and DSS-compatible respectively. We only present results from the gauge choice that leads us closer to the threshold of collapse for each family. In general terms the best gauge choice is one that allows to avoid coordinate singularities and lets us tune further. However, it is worth mentioning that the differences between the cleanest results from each gauge are minor. One of the families is ingoing, denoted as “ing”, whilst the rest start at a moment of time symmetry, denoted as “ts”. The last half-echo of every family is sensitive to numerical error. The more it resembles the earlier peaks, the better the results are. In the case of the ingoing family we see a slightly better last half-echo when using the HDWG-$\ln\left(\alpha\right)$ gauge rather than the DSS-log gauge and this gauge furthermore allowed us to tune two more digits. For these reasons we chose to present this gauge for that particular family.
@@
fitting the values by regression. We find values around $\gamma\simeq 0.37$ for all families. This coincides with the familiar value from simulations of the real scalar field in spherical symmetry [1].
@@ V.2 Departures from Spherical Symmetry
Numerical data produced with bamps has been thoroughly convergence tested in numerous challenging validation runs, see for instance [25, 26, 33]. Doing so carefully is a technically tricky task with any pseudospectral method, because one cannot expect a simple power-law improvement in errors with resolution as in a finite-differencing method. Instead, in the best case scenario one should obtain exponential convergence with resolution, but typically with an unknown exponent. (To avoid misinterpretation: exponential convergence is highly desirable, but complicates quality evaluation). These difficulties are amplified yet further when using mesh-refinement, because as resolution is increased, the grid hierarchies will only match if carefully specified at the outset, which requires more book-keeping to set up. For applications in critical collapse, the situation is yet worse because, by design, we consider spacetimes which depend sensitively on the chosen initial data. Consequently numerical error can have an outsized effect on the outcome of the evolution and can, for instance, even kick the end-state of the evolution from dispersion to black hole formation, and thus end up with quite different levels of constraint violation. For these reasons, and the fact that to do so would be computationally highly expensive, we have not performed systematic convergence tests on all of the data presented here. In particular, our mesh-refinement infrastructure was tested in detail in [26], including with pre-prescribed grid hierarchy. In [33] detailed validation runs were presented with exactly the complex scalar field code we employ presently. For this project specifically we performed such tests for strong data far from the threshold of collapse and saw convincing convergence with resolution. Confidence in the quality of our aspherical near-threshold evolutions is bolstered by a number of facts. First, we monitor constraint violation throughout. The constraints do grow as the dynamics become increasingly violent at the threshold of collapse, but they remain at least five orders of magnitude, and often more, smaller than any curvature scalars with equivalent units that we consider. Second, we have already compared above our spherical runs with the (admittedly limited) available literature for the complex scalar field, and find excellent agreement. Below we will reveal that our results agree qualitatively with published work for the analogous setup for the real scalar field. All these points considered, we are therefore confident in the qualitative behavior we find, especially when considering small asphericities, but do not feel justified in stating strict error bounds, which are left for future work. To do so we will need to perform detailed convergence tests, which is complicated because of the sensitive dependence of our evolutions on initial data, the use of AMR, which is essential in this context, and our pseudospectral approximation, for which the exact rate of decay of error with resolution is hard to establish a priori.
@@
For instance, it is possible for the bifurcation to occur such that subsequent curvature peaks occur ever closer to the origin so that in the limit of infinite tuning, the curvature singularity forms at the origin. On the other hand, curvature singularities could form away from the origin. In either case, the blow-up could exhibit DSS or not. At the level of tuning currently possible in vacuum collapse, evidence suggests that there are families of initial data in which the singularity forms at the origin and others in which it does not. In either case, there is no evidence for strict DSS as in the spherical scalar field setting. But there are families that exhibit approximate DSS. Since the dynamics of the complex scalar field necessarily include those of the vacuum and real scalar field cases, all of this is highly relevant. Yet another possibility, which we do not presently have the tools to diagnose, is a true violation of weak cosmic censorship by the formation of a naked singularity on an open set of parameters $\left(p_{\textrm{min}},p_{\textrm{max}}\right)$ within a family.
@@
Figure 8: In this plot we show a quantitative diagnostic of the degree of asphericity of our data for literature comparison purposes. The quantity being plotted is the maximum difference between the values of the complex field amplitude on the $z-$axis and on the $x-$axis. More precisely, we identify the null coordinates on each axis, $\lambda_{x}$ and $\lambda_{z}$, and we take the differences at points that have the same value of that coordinate, $\lambda_{x}=\lambda_{z}$ The plot reveals the maximum value of this difference against slow time and indicates how aspherical each dataset is at the level of initial data as well as at late times. The spherical family is, by definition, represented by the $x-$axis in this plot. It seems that the first two aspherical families present fluctuations of order $0.01-0.05$ of the field difference between the two axes, but converge to the spherical behavior at late times, where the threshold regime lies. The last aspherical family clearly presents a different pattern than the rest, since the degree of asphericity seems to grow with slow time and peak at $0.4$ after the bifurcation has occurred.
Since the asphericity of the spacetime seems to be the key factor behind the departure from the spherical picture of critical collapse, it is desirable to have an unambiguous measure of the asphericity beyond the initial slice. Even if different codes evolve the same initial data, with $\epsilon$ or $s_{z}$ as a measure of initial asphericity, different numerical implementations and especially foliations might lead to different degrees of apparent asphericity in the time development, which might lead to apparent discrepancies between different studies. To compare with [22, 16] and the very recent single-null work of [49, 17], we assess the maximum difference of the scalar field amplitude at the poles and on the equatorial plane measured in our single-null DSS-adapted coordinates. The result is given in Fig. 8. This is a good option because all published data can be evaluated in such coordinates to make a fair comparison. As we could expect from the fact that we only add an aspherical perturbation to the real part of the solution and not to the imaginary part, our $\epsilon=0.75$ is indeed less aspherical than that investigated in [16]. Interestingly, our $\epsilon=0.99$ family is more aspherical than the $\epsilon=0.75$ data of [17], which could also account for the fact that no bifurcation is observed for that data. See [17] for a detailed comparison and discussion of this.
@@
Taken together, Figs. 7 and 10 indicate that near the threshold, our most aspherical family does not exhibit discrete self-similarity centered at the origin and is thus not a single copy of the Choptuik solution. That said, we could still reasonably describe a portion of the spacetime as approximately DSS about the center. They also show that the gravitational wave component of the spacetime becomes much more significant in the highly aspherical regime. Figure 9 corroborates this, and furthermore suggests that a lack of DSS may be more clearly diagnosed by looking at a more complete set of scalar fields.
@@ VI Conclusions
Our first set of findings concerns spherical symmetry configurations, for which we tuned four distinct families of initial data to the threshold of collapse. Although closely related models have been studied before, as far as we are aware, this is the first time that the spherical configuration for the massless complex field has been carefully treated. We found strong evidence that there is a universal critical spacetime metric which agrees with that of the real scalar field. This metric is characterized partially by the echoing period $\Delta\simeq 3.4$. Perturbations from the critical solution lead to power-law growth of scalar quantities with exponent $\gamma\simeq 0.37$ also familiar from the spherical real scalar field [1]. Since the complex scalar solution space includes that of the real field, this was certainly a possibility a priori, but requires the real and imaginary parts of the field to concoct precisely the stress-energy tensor of the Choptuik solution, and so was a surprise to us. Interestingly similar behavior in which, in a certain limit, the dynamics of a more complicated system agree with those of a simpler model, occurs at the threshold of blow-up for several of the models examined in [53]. This surprise is heightened by the fact that, in the general axisymmetric setting with twist (angular momentum) [22] the threshold of collapse for real and complex fields certainly do not coincide; in that context the real field cannot support angular momentum whereas the complex field can.
@@ Acknowledgments
We acknowledge financial support provided by FCT/Portugal through grants 2022.01324.PTDC, PTDC/FIS-AST/7002/2020, UIDB/00099/2020 and UIDB/04459/2020. D.C. acknowledges support from the by the STFC Research Grant No. ST/V005669/1 “Probing Fundamental Physics with Gravitational-Wave Observations”. D.C. and F.A. were supported in part by the Deutsche Forschungsgemeinschaft (DFG) under Grant No. 406116891 within the Research Training Group RTG 2522/1. H.R.R acknowledges support from the Fundação para a Ciência e Tecnologia (FCT) within the projects UID/04564/2021, UIDB/04564/2020, UIDP/04564/2020 and EXPL/FIS-AST/0735/2021 and the H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. The figures in this article were produced with Plots.jl [70], Matplotlib [71], A. Khirnov’s nr_analysis_axi package [72] and ParaView [73, 74]. Several calculations were performed using the xAct package [75] for Mathematica. Numerical simulations were performed at the Leibniz Supercomputing Centre (LRZ), supported by the project pn36je. The authors thankfully acknowledge the computer resources, technical expertise and assistance provided by CENTRA/IST. Computations were performed at the cluster “Baltasar-Sete-Sóis” and supported by the H2020 ERC Consolidator Grant “Matter and strong field gravity: New frontiers in Einstein’s theory” grant agreement no. MaGRaTh-646597.